Travis purchased a house for $325,000. He made a down payment of 20.00% of the value of the house and received a mortgage for the rest of the amount at 3.42% compounded semi-annually amortized over 15 years. The interest rate was fixed for a 6 year period.

Calculate the monthly payment amount.

Calculate the principal balance at the end of the 6 year term.

Calculate the monthly payment amount if the mortgage was renewed for another 6 years at 4.02% compounded semi-annually?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

Purchase Price of House Down Payment Loan amount Term in years Term in months $325,000.00 20% $260,000.00 180 For, first 6 ye

Cell reference -

B Purchase Price of House Down Payment Loan amount Term in years Term in months 325000 0.2 =C2*(1-C3) 15 =C5*12 6 8 9 10 11 a

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.